OpenStudy (anonymous):

2square root of 6 * squareroot of -3 would this be no solution?

OpenStudy (anonymous):

u can take a square root of negative??, im sure u cant..

OpenStudy (phi):

are you studying complex numbers or imaginary numbers?

OpenStudy (anonymous):

no regulars we havnt done imaginary numbers

OpenStudy (xishem):

\[2 \sqrt{6}*\sqrt{-3}=2 \sqrt{6}*i \sqrt{3}=6i \sqrt{2}\]

OpenStudy (xishem):

Ignore that then.

OpenStudy (anonymous):

then it would be no solution since you didn't learn imaginary numbers yet